Mary is a smart entrepreneur in her primary school. She sells ball points to her fellow school mates. At a price of $2.00 each, she sells 100. At a price of $1.50 each, she sells 300. 

    (i)  Is Mary’s demand elastic or inelastic over this price range?       

    (ii) If demand had the same elasticity for a price decline from $1.50 to $1.00 as it does for the decline from $2.00 to $1.50, would cutting the price from $1.50 to $1.00 increase or decrease Mary’s total revenue?     



1
Expert's answer
2021-01-30T06:30:49-0500

(i) Price elasticity of demand is:

"Ed = \frac{300 - 100} {1.5 - 3} ×\frac{1.5 + 3} {300 + 100} = -1.5."

So, Mary’s demand is elastic over this price range.

(ii) If demand had the same elasticity for a price decline from $1.50 to $1.00 as it does for the decline from $2.00 to $1.50, then cutting the price from $1.50 to $1.00 would increase Mary’s total revenue.